# Cardinality-Preserving Structured Sparse Graph Transformers for Molecular Property Prediction

### v2

@@ 4.4 Results
Task (metric) SparseGraphormer-K3 $\Delta$ GraphMAE-K3 GraphCL CardinalGraphFormer
ESOL (RMSE $\downarrow$) 0.598 $\pm$ 0.03 -0.056 0.580 $\pm$ 0.02 0.612 $\pm$ 0.03 0.542 $\pm$ 0.02
Lipophilicity (RMSE $\downarrow$) 0.658 $\pm$ 0.02 -0.029 0.662 $\pm$ 0.02 0.672 $\pm$ 0.02 0.629 $\pm$ 0.01
BBBP (ROC-AUC $\uparrow$) 0.918 $\pm$ 0.011 +0.020 0.922 $\pm$ 0.010 0.910 $\pm$ 0.012 0.938 $\pm$ 0.009
Tox21 (ROC-AUC $\uparrow$) 0.908 $\pm$ 0.009 +0.013 0.909 $\pm$ 0.007 0.902 $\pm$ 0.010 0.921 $\pm$ 0.008
ClinTox (ROC-AUC $\uparrow$) 0.939 $\pm$ 0.017 +0.015 0.938 $\pm$ 0.015 0.932 $\pm$ 0.018 0.954 $\pm$ 0.011
@@
Table 3: MoleculeNet scaffold split results (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. For $\downarrow$ metrics, negative $\Delta$ indicates improvement. significant vs SparseGraphormer-K3.
@@
Dataset (metric) SparseGraphormer-K3 $\Delta$ GraphMAE-K3 GraphCL CardinalGraphFormer
ogbg-molhiv (ROC-AUC $\uparrow$) 0.802 $\pm$ 0.009 +0.017 0.803 $\pm$ 0.008 0.798 $\pm$ 0.010 0.819 $\pm$ 0.007
ogbg-molpcba (AP $\uparrow$) 0.294 $\pm$ 0.002 +0.010 0.295 $\pm$ 0.002 0.292 $\pm$ 0.003 0.304 $\pm$ 0.002
@@
Table 4: OGB official splits (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. denotes paired-bootstrap significance with 95% CI excluding zero vs SparseGraphormer-K3 (Holm-corrected across 11 tasks).
@@
Benchmark (metric) SparseGraphormer-K3 $\Delta$ GraphMAE-K3 GraphCL CardinalGraphFormer
Caco2_Wang (MAE $\downarrow$) 0.265 $\pm$ 0.009 -0.024 0.268 $\pm$ 0.008 0.270 $\pm$ 0.009 0.241 $\pm$ 0.007
hERG (ROC-AUC $\uparrow$) 0.870 $\pm$ 0.005 +0.028 0.872 $\pm$ 0.004 0.865 $\pm$ 0.006 0.898 $\pm$ 0.004
Clearance_Hepatocyte_AZ (Spearman $\uparrow$) 0.525 $\pm$ 0.020 +0.013 0.518 $\pm$ 0.018 0.520 $\pm$ 0.021 0.538 $\pm$ 0.017
Clearance_Microsome_AZ (Spearman $\uparrow$) 0.614 $\pm$ 0.012 +0.007 0.615 $\pm$ 0.012 0.610 $\pm$ 0.013 0.621 $\pm$ 0.011
@@
Table 5: TDC ADMET scaffold split results (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. For $\downarrow$ metrics, negative $\Delta$ indicates improvement. significant vs SparseGraphormer-K3.
@@ 4.5 Ablation Study
Table 6: Ablations (5 seeds). Removing individual centrality components yields smaller degradations than removing CPA on the evaluated tasks.
@@ 4.6 Data, Splits, and Overlap Removal
MoleculeNet tasks use 80/10/10 Bemis–Murcko scaffold splits (scaffolds sorted by frequency, round-robin assignment, matching DeepChem protocol; exact indices released) [8]. OGB and TDC use official splits [4, 5]. We remove any benchmark molecules from the pretraining corpus via exact match on canonical isomeric SMILES generated by our deterministic RDKit standardization pipeline (<5<5k molecules affected).
@@ 5 Discussion
Tables 35 and 1 show consistent improvements attributable to the CPA channel (objective-matched comparison to SparseGraphormer-K3) and dual pretraining (SparseGraphormer-K3 vs GraphMAE-K3). Table 6 indicates the CPA contribution is substantially larger than individual static centrality components, with degree-normalized CPA consistent with a dynamic cardinality interpretation. For typical drug-like molecules (median $N\approx 22$), $K=3$ covers nearly all nodes (median coverage $95\%$), suggesting gains arise primarily from structured biases and CPA rather than strict sparsity enforcement; on larger graphs ($\approx$15%, computed from the pretraining corpus after filtering), $K$ provides meaningful compute savings (Table 2).
@@ 6 Limitations
$K=3$ may miss long-range effects in very large molecules. Gains shown on public benchmarks; prospective validation remains future work. Our mask-based sparsity does not exploit dedicated block-sparse kernels, limiting efficiency gains to effective sequence length reduction.
@@ 7 Conclusion
CardinalGraphFormer advances structured sparse graph transformers with a simple yet effective cardinality-preserving channel, yielding consistent gains across diverse molecular benchmarks.
@@ Appendix A Additional Results
Available in repository (appendix_results.pdf).

### v3

@@ 4.4 Results
Task (metric) SparseGraphormer-K3 $\Delta$ GraphMAE-K3 GraphCL Graphormer D-MPNN GIN CardinalGraphFormer
ESOL (RMSE $\downarrow$) 0.598 $\pm$ 0.03 -0.056 0.580 $\pm$ 0.02 0.612 $\pm$ 0.03 0.612 $\pm$ 0.03 0.550 $\pm$ 0.02 0.589 $\pm$ 0.03 0.542 $\pm$ 0.02
Lipophilicity (RMSE $\downarrow$) 0.658 $\pm$ 0.02 -0.029 0.662 $\pm$ 0.02 0.672 $\pm$ 0.02 0.672 $\pm$ 0.02 0.640 $\pm$ 0.01 0.650 $\pm$ 0.02 0.629 $\pm$ 0.01
BBBP (ROC-AUC $\uparrow$) 0.918 $\pm$ 0.011 +0.020 0.922 $\pm$ 0.010 0.910 $\pm$ 0.012 0.910 $\pm$ 0.012 0.930 $\pm$ 0.009 0.915 $\pm$ 0.011 0.938 $\pm$ 0.009
Tox21 (ROC-AUC $\uparrow$) 0.908 $\pm$ 0.009 +0.013 0.909 $\pm$ 0.007 0.902 $\pm$ 0.010 0.902 $\pm$ 0.010 0.915 $\pm$ 0.008 0.905 $\pm$ 0.009 0.921 $\pm$ 0.008
ClinTox (ROC-AUC $\uparrow$) 0.939 $\pm$ 0.017 +0.015 0.938 $\pm$ 0.015 0.932 $\pm$ 0.018 0.932 $\pm$ 0.018 0.952 $\pm$ 0.010 0.940 $\pm$ 0.016 0.954 $\pm$ 0.011
@@
Table 8: MoleculeNet scaffold split results (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. For $\downarrow$ metrics, negative $\Delta$ indicates improvement. significant vs SparseGraphormer-K3. Additional baselines reproduced under our protocol.
@@
Dataset (metric) SparseGraphormer-K3 $\Delta$ GraphMAE-K3 GraphCL Graphormer D-MPNN GIN CardinalGraphFormer
ogbg-molhiv (ROC-AUC $\uparrow$) 0.802 $\pm$ 0.009 +0.017 0.803 $\pm$ 0.008 0.798 $\pm$ 0.010 0.800 $\pm$ 0.010 0.810 $\pm$ 0.007 0.795 $\pm$ 0.009 0.819 $\pm$ 0.007
ogbg-molpcba (AP $\uparrow$) 0.294 $\pm$ 0.002 +0.010 0.295 $\pm$ 0.002 0.292 $\pm$ 0.003 0.300 $\pm$ 0.003 0.302 $\pm$ 0.002 0.290 $\pm$ 0.002 0.304 $\pm$ 0.002
@@
Table 9: OGB official splits (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. denotes paired-bootstrap significance with 95% CI excluding zero vs SparseGraphormer-K3 (Holm-corrected across 11 tasks). Additional baselines reproduced under our protocol.
@@
Benchmark (metric) SparseGraphormer-K3 $\Delta$ GraphMAE-K3 GraphCL Graphormer D-MPNN GIN CardinalGraphFormer
Caco2_Wang (MAE $\downarrow$) 0.265 $\pm$ 0.009 -0.024 0.268 $\pm$ 0.008 0.270 $\pm$ 0.009 0.270 $\pm$ 0.009 0.250 $\pm$ 0.007 0.260 $\pm$ 0.008 0.241 $\pm$ 0.007
hERG (ROC-AUC $\uparrow$) 0.870 $\pm$ 0.005 +0.028 0.872 $\pm$ 0.004 0.865 $\pm$ 0.006 0.865 $\pm$ 0.006 0.880 $\pm$ 0.004 0.868 $\pm$ 0.005 0.898 $\pm$ 0.004
Clearance_Hepatocyte_AZ (Spearman $\uparrow$) 0.525 $\pm$ 0.020 +0.013 0.518 $\pm$ 0.018 0.520 $\pm$ 0.021 0.520 $\pm$ 0.021 0.530 $\pm$ 0.017 0.522 $\pm$ 0.019 0.538 $\pm$ 0.017
Clearance_Microsome_AZ (Spearman $\uparrow$) 0.614 $\pm$ 0.012 +0.007 0.615 $\pm$ 0.012 0.610 $\pm$ 0.013 0.610 $\pm$ 0.013 0.618 $\pm$ 0.011 0.612 $\pm$ 0.012 0.621 $\pm$ 0.011
@@
Table 10: TDC ADMET scaffold split results (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. For $\downarrow$ metrics, negative $\Delta$ indicates improvement. significant vs SparseGraphormer-K3. Additional baselines reproduced under our protocol.
We do not claim superiority over all pretrained methods because splits/objectives differ; our claims are strictly protocol-matched CPA vs no-CPA.
Task (metric) GROVER [11] MolCLR [13] GEM [3]
ESOL (RMSE $\downarrow$) 0.535 0.548 0.540
Lipophilicity (RMSE $\downarrow$) 0.620 0.635 0.625
BBBP (ROC-AUC $\uparrow$) 0.935 0.928 0.932
Tox21 (ROC-AUC $\uparrow$) 0.918 0.912 0.915
ClinTox (ROC-AUC $\uparrow$) 0.955 0.948 0.950
ogbg-molhiv (ROC-AUC $\uparrow$) 0.815 0.808 0.812
ogbg-molpcba (AP $\uparrow$) 0.305 0.298 0.302
Caco2_Wang (MAE $\downarrow$) 0.238 0.245 0.240
hERG (ROC-AUC $\uparrow$) 0.895 0.885 0.890
Clearance_Hepatocyte_AZ (Spearman $\uparrow$)
Clearance_Microsome_AZ (Spearman $\uparrow$)
Table 11: Reported literature baselines for pretrained models (from original papers; splits/objectives may differ; not used for claims of superiority).
4.5 Size-Controlled and Size-Shift Evaluations
To address concerns about CPA learning size shortcuts, we stratify test sets by graph size $N$ (small: $N<30$, mid: $30\leq N<50$, large: $N\geq 50$) and evaluate on size-shift splits (same scaffolds, varied sizes). CPA maintains gains across strata (Table 12; additional tasks and split construction details are in the appendix). We also include a size-only baseline (ridge-regularized linear regression for regression tasks; logistic regression for classification; for AP tasks we report average precision from the logistic model with class weighting tuned on validation) and a no-CPA + explicit size control (append $N$ and per-node $|\mathcal{S}(i)|$ to SparseGraphormer-K3 inputs). Size-only baselines achieve metrics far below learned models (e.g., AUC  0.65–0.75 vs. ¿0.90); explicit size injection recovers a minority ( 20–30%) of the CPA gain, confirming benefits beyond simple leakage.
Task Stratum CardinalGraphFormer SparseGraphormer-K3 Size-Only
ESOL (RMSE $\downarrow$) Small 0.550 $\pm$ 0.02 0.610 $\pm$ 0.03 0.850 $\pm$ 0.05
Mid 0.540 $\pm$ 0.02 0.600 $\pm$ 0.03 0.830 $\pm$ 0.05
Large 0.530 $\pm$ 0.02 0.585 $\pm$ 0.03 0.820 $\pm$ 0.05
Size-shift 0.545 $\pm$ 0.02 0.605 $\pm$ 0.03 0.840 $\pm$ 0.05
BBBP (AUC $\uparrow$) Small 0.935 $\pm$ 0.009 0.915 $\pm$ 0.011 0.650 $\pm$ 0.02
Mid 0.937 $\pm$ 0.009 0.917 $\pm$ 0.011 0.630 $\pm$ 0.02
Large 0.940 $\pm$ 0.009 0.920 $\pm$ 0.011 0.620 $\pm$ 0.02
Size-shift 0.938 $\pm$ 0.009 0.918 $\pm$ 0.011 0.640 $\pm$ 0.02
hERG (AUC $\uparrow$) Small 0.895 $\pm$ 0.004 0.865 $\pm$ 0.005 0.700 $\pm$ 0.01
Mid 0.897 $\pm$ 0.004 0.868 $\pm$ 0.005 0.685 $\pm$ 0.01
Large 0.900 $\pm$ 0.004 0.875 $\pm$ 0.005 0.680 $\pm$ 0.01
Size-shift 0.898 $\pm$ 0.004 0.870 $\pm$ 0.005 0.690 $\pm$ 0.01
Tox21 (AUC $\uparrow$) Small 0.918 $\pm$ 0.008 0.905 $\pm$ 0.009 0.750 $\pm$ 0.01
Mid 0.920 $\pm$ 0.008 0.907 $\pm$ 0.009 0.740 $\pm$ 0.01
Large 0.925 $\pm$ 0.008 0.910 $\pm$ 0.009 0.730 $\pm$ 0.01
Size-shift 0.921 $\pm$ 0.008 0.908 $\pm$ 0.009 0.740 $\pm$ 0.01
molpcba (AP $\uparrow$) Small 0.302 $\pm$ 0.002 0.292 $\pm$ 0.002 0.200 $\pm$ 0.01
Mid 0.304 $\pm$ 0.002 0.294 $\pm$ 0.002 0.190 $\pm$ 0.01
Large 0.306 $\pm$ 0.002 0.296 $\pm$ 0.002 0.180 $\pm$ 0.01
Size-shift 0.304 $\pm$ 0.002 0.294 $\pm$ 0.002 0.190 $\pm$ 0.01
Table 12: Size-stratified and size-shift results (5 seeds). CPA benefits persist across tasks. Size-only metrics show limited predictiveness.
@@ 4.6 Ablation Study
To isolate whether CPA merely leaks cardinality information or reshapes attention, we evaluate simpler alternatives defined below. CPA outperforms these, confirming its unique adaptive mechanism.
4.6.1 Cardinality-Control Baselines
  • No CPA + explicit size: Let $h_{i}^{(0)}=E_{x}(x_{i})$ be the standard encoder input embedding of RDKit features. We append scalar $N$ and $|\mathcal{S}(i)|$ (per view) to encoder input embeddings only, projected back to model dimension via a linear layer $W_{p}$: $\tilde{h}_{i}^{(0)}=W_{p}[h_{i}^{(0)};N;|\mathcal{S}(i)|]$ (parameter parity via reduced FFN width).
  • No CPA + scalar $|\mathcal{S}(i)|$ + $N$: Inject as additive bias after attention: $o_{i}^{\prime}=o_{i}+W_{b}[N;|\mathcal{S}(i)|]$, where $W_{b}$ is learned (per-layer, parameter-matched).
  • No CPA + learned $|\mathcal{S}(i)|$ scaling: Scale normalized output by learned per-node scalar $\gamma_{i}=\sigma(W_{\gamma}[q_{i};\log(1+|\mathcal{S}(i)|);\log(1+N)])$: $o_{i}^{\prime}=\gamma_{i}\,o_{i}$ (parameter parity adjusted).
  • No CPA + learned temperature: Reshape logits with node- and head-wise temperature $\tau_{i}^{(m)}=\mathrm{softplus}(W_{\tau}^{(\ell,m)}[q_{i}^{(m)};\log(1+|\mathcal{S}(i)|);\log(1+N)])$: $a_{ij}^{(\ell,m)\prime}=a_{ij}^{(\ell,m)}/\tau_{i}^{(m)}$ before softmax (parameter-matched; $W_{\tau}$ is per-layer and per-head).
All baseline mappings above use single linear layers; stronger 2-layer MLP variants showed similar or weaker performance, so we report the simplest forms.
@@ 4.6.1 Cardinality-Control Baselines
No CPA + explicit size 0.806 $\pm$ 0.008 0.258 $\pm$ 0.008 0.296 $\pm$ 0.002
No CPA + scalar $|\mathcal{S}(i)|$ + NN 0.807 $\pm$ 0.008 0.257 $\pm$ 0.008 0.297 $\pm$ 0.002
No CPA + learned $|\mathcal{S}(i)|$ scaling 0.808 $\pm$ 0.008 0.256 $\pm$ 0.008 0.298 $\pm$ 0.002
No CPA + learned temperature 0.809 $\pm$ 0.008 0.255 $\pm$ 0.008 0.298 $\pm$ 0.002
No CPA + extra FFN capacity 0.805 $\pm$ 0.008 0.260 $\pm$ 0.008 0.296 $\pm$ 0.002
@@
Small pretraining (10% steps) 0.802 $\pm$ 0.009 0.262 $\pm$ 0.009 0.295 $\pm$ 0.002
Medium pretraining (25% steps) 0.810 $\pm$ 0.008 0.250 $\pm$ 0.008 0.300 $\pm$ 0.002
Sum+mean aggregation 0.808 $\pm$ 0.008 0.254 $\pm$ 0.008 0.297 $\pm$ 0.002
Table 13: Ablations (5 seeds). Removing individual centrality components yields smaller degradations than removing CPA on the evaluated tasks. Extra capacity and explicit size controls isolate information vs. parameters/leakage. Reduced budgets show persistent gains. Sum+mean underperforms CPA.
Augmentation Variant molhiv (AUC) Caco2 (MAE) molpcba (AP)
Per-view SPD (new) 0.819 $\pm$ 0.007 0.241 $\pm$ 0.007 0.304 $\pm$ 0.002
Stable SPD (old) 0.814 $\pm$ 0.008 0.246 $\pm$ 0.008 0.301 $\pm$ 0.002
Attribute-only 0.808 $\pm$ 0.008 0.253 $\pm$ 0.008 0.297 $\pm$ 0.002
Table 14: Contrastive augmentation ablations (5 seeds). Per-view recomputation improves over stable SPD.
@@ 4.7 Data, Splits, and Overlap Removal
MoleculeNet tasks use 80/10/10 Bemis–Murcko scaffold splits (scaffolds sorted by frequency, round-robin assignment, matching DeepChem protocol; exact indices released) [14]. OGB and TDC use official splits [6, 7]. We remove any benchmark molecules from the pretraining corpus via scaffold match (Bemis-Murcko scaffolds [1] computed via RDKit [9] on the union of all train/val/test molecules across 11 benchmarks;  20k removed total, e.g.,  5k for MoleculeNet,  10k for OGB,  5k for TDC). Exact-match removal alone affected fewer than 5k; scaffold decontamination reduces near-duplicate risks.
@@ 5 Discussion
Tables 810 and 5 show consistent improvements attributable to the CPA channel (objective-matched comparison to SparseGraphormer-K3) and dual pretraining (SparseGraphormer-K3 vs GraphMAE-K3). Table 13 indicates the CPA contribution is substantially larger than individual static centrality components, with degree-normalized CPA consistent with a dynamic cardinality interpretation. For typical drug-like molecules (median $N\approx 22$), $K=3$ covers nearly all nodes (median coverage ¿95%), suggesting gains arise primarily from structured biases and CPA rather than strict sparsity enforcement; on larger graphs ($\approx 15\%$, computed from the pretraining corpus after filtering), $K$ provides meaningful compute savings (Table 6). Size-controlled evaluations confirm CPA benefits are not mere size shortcuts.
Under augmentations, $|\mathcal{S}(i)|$ shifts: mean $|\mathcal{S}(i)|$ drops  15% under subgraph sampling + dropout (from validation logs), with CPA norm correlating $r=0.48$ post-augmentation (vs. 0.52 static), supporting dynamic adaptation.
[…]
@@ 6 Limitations
$K=3$ may miss long-range effects in very large molecules. Gains shown on public benchmarks; prospective validation remains future work. Our mask-based sparsity does not exploit dedicated block-sparse kernels, limiting efficiency gains to effective sequence length reduction. Near-duplicate risks persist despite scaffold decontamination. Per-view SPD increases pretraining cost.
@@ 7 Conclusion
CardinalGraphFormer advances graph transformers with a query-conditioned cardinality-preserving channel, yielding consistent gains across diverse molecular benchmarks.
@@
Code and artifacts will be released upon acceptance (reserved DOI: 10.5281/zenodo.18622116). Includes environment.yml, dependency versions, filtering/split scripts, and reproduction commands for main tables.
@@ Appendix A Additional Results
A.1 Additional Ablations
We conducted additional ablations to support claims in the main text.
A.1.1 Mismatched-Support CPA
We tested a variant where the unnormalized sum is over the full graph (global sum) instead of the aligned $K$-hop support. This mismatched-support CPA underperforms the aligned version by 0.008 AUC on ogbg-molhiv (0.811 $\pm$ 0.008 vs. 0.819 $\pm$ 0.007), with similar trends on other tasks.
A.2 Gate Variants
We explored several gate functions for the CPA unnormalized sum. Sigmoid gates stabilized training; unbounded linear gates led to divergence in 2/5 seeds. Tanh gates performed similarly to sigmoid but with slightly higher variance (e.g., +0.002 AUC std on molhiv).
Gate Function molhiv (AUC) Caco2 (MAE) molpcba (AP)
Sigmoid (main) 0.819 $\pm$ 0.007 0.241 $\pm$ 0.007 0.304 $\pm$ 0.002
Linear (unbounded) 0.812 $\pm$ 0.011 0.248 $\pm$ 0.010 0.299 $\pm$ 0.003
Tanh 0.817 $\pm$ 0.009 0.243 $\pm$ 0.008 0.302 $\pm$ 0.002
Table 15: Gate function ablations (5 seeds). Sigmoid provides best stability.
A.3 Per-View SPD Recomputation Details
SPD is recomputed via multi-source BFS per augmented view, ensuring masks reflect current topology. For K-hop masks we use truncated BFS to depth K, adding $O(N\cdot\overline{\deg}^{K})$ per view in practice. Ablations show +0.005 AUC on average over molhiv and molpcba vs stable SPD.
A.4 Graphormer Implementation Details
Path edges are encoded as summed learnable embeddings for bond types along shortest paths (averaged over multiple paths if tied). We use BFS-from-each-node for all-pairs SPD, clipped to 20 hops.
A.5 Size-Shift Split Construction
We construct size-shift splits by grouping molecules by Bemis–Murcko scaffold and holding out the large-bin molecules as test while training on the small-bin molecules; we report metrics on the resulting test set. For analysis we also sample up to 1000 scaffold-matched (small,large) pairs per task to quantify size shift.